OpenStudy (anonymous):

A shell is fired from a long-range gun with an initial velocity = 300 m/s at 30 degrees on a flat surface. Find the height of the projectile at t=20s. Answer key says 1000 m, can't get the right answer!

OpenStudy (anonymous):

I tried change in x = V0(t) +1/2(a)(t^2) x=300(20) + 1/2(-10)(20^2) = 4000 m, but doesn't match up to 1000m (We use 10 as a simplified g)

OpenStudy (wolfe8):

The bullet was shot at an angle, so you will have to resolve the velocity into x and y components.

OpenStudy (wolfe8):



OpenStudy (anonymous):

So Vx = 300cos30, Vy = 300sin30

OpenStudy (anonymous):

I was trying to find formulas with t so I can substitute it in

OpenStudy (wolfe8):

Yup. So use the formula for range. I can't remember which one it is sorry. I'm exhausted.

OpenStudy (anonymous):

Ah ok. Will try. Thanks and gl on ur exam!

OpenStudy (anonymous):

Does anyone know the right formula? I'm confused which one to use and how to set it up

OpenStudy (anonymous):

Got it, it is Y = \[Vyi (t) - 1/2 (g)(t^2) = 300\sin30(2) - 1/2 (10)(20^2)\] = 3000-2000 = 1000m , using g = 10